# BOP Challenge 2023 on Detection, Segmentation and Pose Estimation of Seen and Unseen Rigid Objects

### v1

@@ Abstract
We present the evaluation methodology, datasets and results of the BOP Challenge 2023, the fifth in a series of public competitions organized to capture the state of the art in model-based 6D object pose estimation from an RGB/RGB-D image and related tasks. Besides the three tasks from 2022 (2D detection, 2D segmentation, and 6D localization of objects seen during training), the 2023 challenge introduced new variants of these tasks focused on objects unseen during training. In the new tasks, methods were required to learn new objects during a short onboarding stage (max 5 minutes, 1 GPU) from provided 3D object models. The best 2023 method for 6D localization of unseen objects (GenFlow) notably reached the accuracy of the best 2020 method for seen objects (CosyPose), although being noticeably slower. The best 2023 method for seen objects (GPose) achieved a moderate accuracy improvement but a significant 43% run-time improvement compared to the best 2022 counterpart (GDRNPP). Since 2017, the accuracy of 6D localization of seen objects has improved by more than 50% (from 56.9 to 85.6 AR). The online evaluation system stays open and is available at: bop.felk.cvut.cz.
@@ 1 Introduction
Participants of the 2023 challenge were competing on six tasks. Besides the three tasks from 2022 (model-based 2D object detection, 2D object segmentation and 6D object localization of objects seen during training), the 2023 challenge introduced new variants of these tasks focused on objects unseen during training. In the new tasks, methods were required to adapt to novel 3D object models during a short object onboarding stage (max 5 min per object, 1 GPU), and then recognize the objects in images from diverse environments. Such methods are of a high practical relevance as they do not require expensive data generation and training for every new object, which is typically required by most existing methods for seen objects and severely limits their scalability. The introduction of the new tasks was encouraged by the recent breakthroughs in foundation models and their impressive few-shot learning capabilities.
@@
The best 2023 method for 6D localization of unseen objects (GenFlow [41]) reached the accuracy of the best 2020 method for seen objects (CosyPose [33]). Despite being noticeably slower, this is an impressive result considering that the target objects are onboarded in a short time, which is several orders of magnitude shorter than a typical training process of methods trained for specific objects. The best 2023 method for seen objects (GPose [1]) achieves a moderate accuracy improvement and a significant 42.6% run-time improvement compared to the best 2022 counterpart (GDRNPP [62, 40]).
@@ 2.1 Task 1: 6D localization of seen objects
The definition of this task is the same since 2019, which enables direct comparison across the years.333See Sec. A.1 in [26] for a discussion on why the methods are evaluated on 6D object localization instead of 6D object detection, where no prior information about the visible object instances is provided [24].
@@
Evaluation methodology: The error of an estimated pose w.r.t. the ground-truth pose is calculated by three pose-error functions (see Section 2.2 of [26] for details): (1) VSD (Visible Surface Discrepancy) which treats indistinguishable poses as equivalent by considering only the visible object part, (2) MSSD (Maximum Symmetry-Aware Surface Distance) which considers a set of pre-identified global object symmetries and measures the surface deviation in 3D, (3) MSPD (Maximum Symmetry-Aware Projection Distance) which considers the object symmetries and measures the perceivable deviation.
An estimated pose is considered correct w.r.t. a pose-error function , if , where and is the threshold of correctness. The fraction of annotated object instances for which a correct pose is estimated is referred to as Recall. The Average Recall w.r.t. a function , denoted as , is defined as the average of the Recall rates calculated for multiple settings of the threshold and also for multiple settings of a misalignment tolerance in the case of VSD. The accuracy of a method on a dataset is measured by: , which is calculated over estimated poses of all objects from . The overall accuracy on the core datasets is measured by defined as the average of the per-dataset scores (see Section 2.4 of [26] for details).444When calculating AR, scores are not averaged over objects before averaging over datasets, which is done when calculating (Sec. 2.2) to comply with the original COCO evaluation methodology [37].
@@ 2.2 Task 2: 2D detection of seen objects
Evaluation methodology: Following the the evaluation methodology from the COCO 2020 Object Detection Challenge [37], the detection accuracy is measured by the Average Precision (AP). Specifically, a per-object score is calculated by averaging the precision at multiple Intersection over Union (IoU) thresholds: . The accuracy of a method on a dataset is measured by calculated by averaging per-object scores, and the overall accuracy on the core datasets (Sec. 3) is measured by defined as the average of the per-dataset scores. Analagous to the 6D localization task, only annotated object instances for which at least of the projected surface area is visible need to be detected. Correct predictions for instances that are visible from less than are filtered out and not counted as false positives. Up to predictions per image with the highest confidences are considered.
@@ 2.3 Task 3: 2D segmentation of seen objects
Training input: At training time, a method is provided a set of RGB-D training images showing objects that are annotated with ground-truth 2D binary masks. The masks are modal, i.e., covering only the visible object part. The method can also use 3D mesh models that are available for the objects.
@@ 2.4 Task 4: 6D localization of unseen objects
Object-onboarding input: The method is provided 3D mesh models of test objects that were not seen during training. To onboard each object (e.g. to render images/templates or fine-tune a neural network), the method can spend up to 5 minutes of the wall-clock time on a computer with a single GPU. The time is measured from the point right after the raw data (e.g. 3D mesh models) is loaded to the point when the object is onboarded. The method can render images of the 3D object models but cannot use any real images of the objects for onboarding. The object representation (which may be given by a set of templates, an machine-learning model, etc.) needs to be fixed after onboarding (it cannot be updated on test images).
@@ 2.6 Task 6: 2D segmentation of unseen objects
Training input: At training time, a method is provided a set of RGB-D training images showing training objects that are annotated with ground-truth 2D binary masks. The masks are modal, i.e., covering only the visible object part. The method can also use 3D mesh models that are available for the objects.
@@
| LM-O [2] | 8 |
50K |
1214 |
200 |
9038 |
1445 |
| T-LESS [23] | 30 |
37584 |
50K |
10080 |
1000 |
67308 |
6423 |
| ITODD [10] | 28 |
50K |
54 |
721 |
721 |
3041 |
3041 |
| HB [30] | 33 |
50K |
4420 |
13000 |
300 |
67542 |
1630 |
| YCB-V [65] | 21 |
113198 |
50K |
20738 |
900 |
98547 |
4123 |
| TUD-L [25] | 3 |
38288 |
50K |
23914 |
600 |
23914 |
600 |
| IC-BIN [8] | 2 |
50K |
177 |
150 |
2176 |
1786 |
| LM [18] | 15 |
50K |
18273 |
3000 |
18273 |
3000 |
| RU-APC [51] | 14 |
5964 |
1380 |
5964 |
1380 |
| IC-MI [58] | 6 |
2067 |
300 |
5318 |
800 |
| TYO-L [25] | 21 |
1670 |
1670 |
1670 |
1670 |
| HOPE [59] | 28 |
50 |
188 |
188 |
3472 |
2898 |
@@ 3.1 Core datasets
BOP currently includes twelve datasets in a unified format. Sample test images are in Fig. 3 and dataset parameters in Tab. 1. Seven from the twelve were selected as core datasets: LM-O, T-LESS, ITODD, HB, YCB-V, TUD-L, IC-BIN. Since 2019, methods must be evaluated on all of these core datasets to be considered for the main challenge awards (Sec. 5).
@@ 3.2 Training dataset for tasks on unseen objects
In 2023, as a training dataset for Tasks 4–6 (Sec. 2), we provided over 2M images in the BOP format showing more than 50K diverse objects (Fig. 3). The images were originally synthesized for MegaPose [34] using BlenderProc [6, 5, 7]. The objects are from the Google Scanned Objects [9] and ShapeNetCore [3] datasets. Note that symmetry transformations are not available for these objects, but could be identified as described in Section 2.3 of [26].
@@ 4.1 Experimental setup
Participants of the 2023 challenge were submitting results to the online evaluation system at bop.felk.cvut.cz from June 7, 2023 until the deadline on September 28, 2023. The evaluation scripts are publicly available in the BOP toolkit.555github.com/thodan/bop_toolkit
@@
Not a single pixel of test images may have been used for training and onboarding, nor the individual ground-truth annotations that are publicly available for test images of some datasets. Ranges of the azimuth and elevation camera angles, and a range of the camera-object distances determined by the ground-truth poses from test images is the only information about the test set that may have been used during training and onboarding. Only subsets of test images were used (see Tab. 1) to remove redundancies and speed up the evaluation, and only object instances for which at least of the projected surface area is visible were considered in the evaluation.
| # | Year | Type | DNN per | Det./seg. | Refinement | Train im. | …type | Test im. | Method |
LM-O |
T-LESS |
TUD-L |
IC-BIN |
ITODD |
HB |
YCB-V |
Time |
| 1 | GPose2023 [62, 40] | 2023 | DNN | Object | Custom | Coord-guided | RGB-D | PBR+real | RGB-D | 79.4 | 91.4 | 96.4 | 73.7 | 70.4 | 95.0 | 92.8 | 85.6 |
2.67 |
| 2 | GPose2023-OfficialDet [62, 40] | 2023 | DNN | Object | Default GDRNPPDet | Coord-guided | RGB-D | PBR+real | RGB-D | 80.5 | 89.5 | 96.6 | 73.4 | 68.7 | 94.4 | 92.9 | 85.1 |
4.57 |
| 3 | GPose2023-PBR [62, 40] | 2023 | DNN | Object | Custom | Coord-guided | RGB-D | PBR+real | RGB-D | 79.4 | 89.0 | 93.1 | 73.7 | 70.4 | 95.0 | 90.1 | 84.4 |
2.86 |
| 4 | GDRNPP-PBRReal-RGBD-MModel [62, 40] | 2022 | DNN | Object | YOLOX | CIR | RGB-D | PBR+real | RGB-D | 77.5 | 87.4 | 96.6 | 72.2 | 67.9 | 92.6 | 92.1 | 83.7 |
6.26 |
| 5 | GDRNPP-PBR-RGBD-MModel [62, 40] | 2022 | DNN | Object | YOLOX | CIR | RGB-D | PBR | RGB-D | 77.5 | 85.2 | 92.9 | 72.2 | 67.9 | 92.6 | 90.6 | 82.7 |
6.26 |
| 6 | ZebraPoseSAT-EffnetB4-refined [54] | 2023 | DNN | Object | Default GDRNPPDet | CIR | RGB-D | PBR+real | RGB-D | 78.0 | 86.2 | 95.6 | 65.4 | 61.8 | 92.1 | 89.9 | 81.3 |
2.57 |
| 7 | GDRNPP-PBRReal-RGBD-MModel-Fast [62, 40] | 2022 | DNN | Object | YOLOX | Depth adjust. | RGB | PBR+real | RGB-D | 79.2 | 87.2 | 93.6 | 70.2 | 58.8 | 90.9 | 83.4 | 80.5 |
0.23 |
| 8 | OfficialDet-PFA-Mixpbr-RGB-D [28] | 2023 | DNN | Dataset | Default GDRNPPDet | PFA | RGB | PBR+real | RGB-D | 79.2 | 84.9 | 96.3 | 70.6 | 52.6 | 86.7 | 89.9 | 80.0 |
1.19 |
| 9 | GDRNPP-PBRReal-RGBD-MModel-Offi. [62, 40] | 2022 | DNN | Object | Default (synt+real) | CIR | RGB-D | PBR+real | RGB-D | 75.8 | 82.4 | 96.6 | 70.8 | 54.3 | 89.0 | 89.6 | 79.8 |
6.41 |
@@
| 11 | Extended_FCOS+PFA-MixPBR-RGBD [28] | 2022 | DNN | Dataset | Extended FCOS | PFA | RGB | PBR+real | RGB-D | 79.7 | 85.0 | 96.0 | 67.6 | 46.9 | 86.9 | 88.8 | 78.7 |
2.32 |
| 12 | Extended_FCOS+PFA-MixPBR-RGBD-Fast [28] | 2022 | DNN | Dataset | Extended FCOS | PFA | RGB | PBR+real | RGB-D | 79.2 | 77.9 | 95.8 | 67.1 | 46.0 | 86.0 | 88.0 | 77.1 |
0.64 |
| 13 | RCVPose3D-SingleModel-VIVO-PBR [64] | 2022 | DNN | Dataset | RCVPose3D | ICP | RGB-D | PBR+real | RGB-D | 72.9 | 70.8 | 96.6 | 73.3 | 53.6 | 86.3 | 84.3 | 76.8 |
1.34 |
| 14 | ZebraPoseSAT-EffnetB4+ICP(DefaultDet) [54] | 2022 | DNN | Object | Default (synt+real) | ICP | RGB | PBR+real | RGB-D | 75.2 | 72.7 | 94.8 | 65.2 | 52.7 | 88.3 | 86.6 | 76.5 |
0.50 |
| 15 | Extended_FCOS+PFA-PBR-RGBD [28] | 2022 | DNN | Dataset | Extended FCOS | PFA | RGB | PBR | RGB-D | 79.7 | 80.2 | 89.3 | 67.6 | 46.9 | 86.9 | 82.6 | 76.2 |
2.63 |
| 16 | SurfEmb-PBR-RGBD [16] | 2022 | DNN | Dataset | Default (PBR) | Custom | RGB-D | PBR | RGB-D | 76.0 | 82.8 | 85.4 | 65.9 | 53.8 | 86.6 | 79.9 | 75.8 |
9.05 |
| 17 | ZebraPoseSAT-EffnetB4 [54] | 2023 | DNN | Object | Default GDRNPPDet | – | RGB | PBR+real | RGB | 72.9 | 82.1 | 85.0 | 59.2 | 50.4 | 92.2 | 82.8 | 74.9 |
2.50 |
| 18 | GDRNPP-PBRReal-RGBD-SModel [62, 40] | 2022 | DNN | Dataset | YOLOX | Depth adjust. | RGB | PBR+real | RGB-D | 75.7 | 85.6 | 90.6 | 68.0 | 35.6 | 86.4 | 81.7 | 74.8 |
0.56 |
@@
| 20 | Coupled Iterative Refinement (CIR) [38] | 2022 | DNN | Dataset | Default (synt+real) | CIR | RGB-D | PBR+real | RGB-D | 73.4 | 77.6 | 96.8 | 67.6 | 38.1 | 75.7 | 89.3 | 74.1 |
| 21 | GPose2023-RGB [62, 40] | 2023 | DNN | Object | GDet2023 | CIR | RGB | PBR | RGB | 69.9 | 79.9 | 83.1 | 62.6 | 46.0 | 87.6 | 80.9 | 72.9 |
0.24 |
| 22 | GDRNPP-PBRReal-RGB-MModel[62, 40] | 2022 | DNN | Object | YOLOX | – | RGB | PBR+real | RGB | 71.3 | 78.6 | 83.1 | 62.3 | 44.8 | 86.9 | 82.5 | 72.8 |
0.23 |
| 23 | ZebraPoseSAT-EffnetB4 [54] | 2022 | DNN | Object | FCOS | – | RGB | PBR+real | RGB | 72.1 | 80.6 | 85.0 | 54.5 | 41.0 | 88.2 | 83.0 | 72.0 |
0.25 |
| 24 | ZebraPoseSAT-EffnetB4(DefaultDet) [54] | 2022 | DNN | Object | Default (synt+real) | – | RGB | PBR+real | RGB | 70.7 | 76.8 | 84.9 | 59.7 | 41.7 | 88.7 | 81.6 | 72.0 |
0.25 |
| 25 | ZebraPoseSAT-EffnetB4(PBR-DefaultDet) [54] | 2023 | DNN | Object | Default GDRNPPDet | – | RGB | PBR+real | RGB | 72.9 | 81.1 | 75.6 | 59.2 | 50.4 | 92.1 | 72.9 | 72.0 |
0.25 |
| 26 | ZebraPose-SAT [54] | 2022 | DNN | Object | FCOS | – | RGB | PBR+real | RGB | 72.1 | 78.7 | 86.1 | 54.9 | 37.9 | 84.7 | 82.8 | 71.0 |
| 27 | Extended_FCOS+PFA-MixPBR-RGB [28] | 2022 | DNN | Dataset | Extended FCOS | PFA | RGB | PBR+real | RGB | 74.5 | 77.8 | 83.9 | 60.0 | 35.3 | 84.1 | 80.6 | 70.9 |
3.02 |
| 28 | GDRNPP-PBR-RGB-MModel [62, 40] | 2022 | DNN | Object | YOLOX | – | RGB | PBR | RGB | 71.3 | 79.6 | 75.2 | 62.3 | 44.8 | 86.9 | 71.3 | 70.2 |
0.28 |
@@
| 31 | MRPE-PBRReal-RGB-SModel | 2023 | DNN | Dataset | Default GDRNPPDet | Render & com. | RGB | PBR+real | RGB-D | 74.4 | 75.8 | 82.4 | 55.0 | 36.8 | 77.0 | 84.3 | 69.4 |
0.10 |
| 32 | GDRNPP-PBRReal-RGB-SModel | 2022 | DNN | Dataset | YOLOX | CIR | RGB | PBR+real | RGB | 68.6 | 77.6 | 82.7 | 61.7 | 26.0 | 80.9 | 76.8 | 67.8 |
0.46 |

### v2

@@ Abstract
We present the evaluation methodology, datasets and results of the BOP Challenge 2023, the fifth in a series of public competitions organized to capture the state of the art in model-based 6D object pose estimation from an RGB/RGB-D image and related tasks. Besides the three tasks from 2022 (2D detection, 2D segmentation, and 6D localization of objects seen during training), the 2023 challenge introduced new variants of these tasks focused on objects unseen during training. In the new tasks, methods were required to learn new objects during a short onboarding stage (max 5 minutes, 1 GPU) from provided 3D object models. The best 2023 method for 6D localization of unseen objects (GenFlow) notably reached the accuracy of the best 2020 method for seen objects (CosyPose), although being noticeably slower. The best 2023 method for seen objects (GPose) achieved a moderate accuracy improvement but a significant 43% run time improvement compared to the best 2022 counterpart (GDRNPP). Since 2017, the accuracy of 6D localization of seen objects has improved by more than 50% (from 56.9 to 85.6 ARC). The online evaluation system stays open and is available at: bop.felk.cvut.cz.
@@ 1 Introduction
Participants of the 2023 challenge were competing on six tasks. Besides the three tasks from 2022 (model-based 2D object detection, 2D object segmentation and 6D object localization of objects seen during training), the 2023 challenge introduced new variants of these tasks focused on objects unseen during training. In the new tasks, methods were required to adapt to novel 3D object models during a short object onboarding stage (max 5 min per object, 1 GPU), and then recognize the objects in images from diverse environments. Such methods are of high practical relevance as they do not require expensive data generation and training for every new object, which is typically required by most existing methods for seen objects and severely limits their scalability. The introduction of the new tasks was encouraged by the recent breakthroughs in foundation models and their impressive few-shot learning capabilities.
@@
The best 2023 method for 6D localization of unseen objects (GenFlow [40]) reached the accuracy of the best 2020 method for seen objects (CosyPose [32]). Despite being noticeably slower, this is an impressive result considering that the target objects are onboarded in a short time, which is several orders of magnitude shorter than a typical training process of methods trained for specific objects. The best 2023 method for seen objects (GPose [67]) achieves a moderate accuracy improvement and a significant 42.6% run time improvement compared to the best 2022 counterpart (GDRNPP [61, 39]).
@@ 2.1 Task 1: 6D localization of seen objects
The definition of this task is the same since 2019, which enables direct comparison across the years333See Sec. A.1 in [25] for a discussion on why the methods are evaluated on 6D object localization instead of 6D object detection, where no prior information about the visible object instances is provided [23]..
@@
Evaluation methodology: The error of an estimated pose w.r.t. the ground-truth pose is calculated by three pose-error functions (see Sec. 2.2 of [25] for details): (1) VSD (Visible Surface Discrepancy) which treats indistinguishable poses as equivalent by considering only the visible object part, (2) MSSD (Maximum Symmetry-Aware Surface Distance) which considers a set of pre-identified global object symmetries and measures the surface deviation in 3D, (3) MSPD (Maximum Symmetry-Aware Projection Distance) which considers the object symmetries and measures the perceivable deviation.
An estimated pose is considered correct w.r.t. a pose-error function , if , where and is the threshold of correctness. The fraction of annotated object instances for which a correct pose is estimated is referred to as Recall. The Average Recall w.r.t. a function , denoted as , is defined as the average of the Recall rates calculated for multiple settings of the threshold and also for multiple settings of a misalignment tolerance in the case of VSD. The accuracy of a method on a dataset is measured by: , which is calculated over estimated poses of all objects from . The overall accuracy on the core datasets is measured by defined as the average of the per-dataset scores (see Sec. 2.4 of [25] for details)444When calculating ARC, scores are not averaged over objects before averaging over datasets, which is done when calculating (Sec. 2.2) to comply with the original COCO evaluation methodology [36]..
@@ 2.2 Task 2: 2D detection of seen objects
Test input: At test time, the method is given an RGB-D image unseen during training that shows an arbitrary number of instances of an arbitrary number of objects, with all objects being from one specified dataset (e.g. YCB-V [64]). No prior information about the visible object instances is provided.
@@
Evaluation methodology: Following the evaluation methodology from the COCO 2020 Object Detection Challenge [36], the detection accuracy is measured by the Average Precision (AP). Specifically, a per-object score is calculated by averaging the precision at multiple Intersection over Union (IoU) thresholds: . The accuracy of a method on a dataset is measured by calculated by averaging per-object scores, and the overall accuracy on the core datasets (Sec. 3) is measured by defined as the average of the per-dataset scores. Analogous to the 6D localization task, only annotated object instances for which at least of the projected surface area is visible need to be detected. Correct predictions for instances that are visible from less than are filtered out and not counted as false positives. Up to predictions per image with the highest confidences are considered.
@@ 2.3 Task 3: 2D segmentation of seen objects
Training input: At training time, a method is provided a set of RGB-D training images showing objects that are annotated with ground-truth 2D binary masks. The masks are modal, i.e., covering only the visible object parts. The method can also use 3D mesh models that are available for the objects.
@@ 2.4 Task 4: 6D localization of unseen objects
Object-onboarding input: The method is provided 3D mesh models of test objects that were not seen during training. To onboard each object (e.g. to render images/templates or fine-tune a neural network), the method can spend up to 5 minutes of the wall-clock time on a computer with a single GPU. The time is measured from the point right after the raw data (e.g. 3D mesh models) is loaded to the point when the object is onboarded. The method can render images of the 3D object models but cannot use any real images of the objects for onboarding. The object representation (which may be given by a set of templates, a machine-learning model, etc.) needs to be fixed after onboarding (it cannot be updated on test images).
@@ 2.6 Task 6: 2D segmentation of unseen objects
Training input: At training time, a method is provided a set of RGB-D training images showing training objects that are annotated with ground-truth 2D binary masks. The masks are modal, i.e., covering only the visible object parts. The method can also use 3D mesh models that are available for the training objects.
@@
| LM-O [1] | 8 | – | 50K | – | 1214 | 200 | 9038 | 1445 |
| T-LESS [22] | 30 | 37584 | 50K | – | 10080 | 1000 | 67308 | 6423 |
| ITODD [9] | 28 | – | 50K | 54 | 721 | 721 | 3041 | 3041 |
| HB [29] | 33 | – | 50K | 4420 | 13000 | 300 | 67542 | 1630 |
| YCB-V [64] | 21 | 113198 | 50K | – | 20738 | 900 | 98547 | 4123 |
| TUD-L [24] | 3 | 38288 | 50K | – | 23914 | 600 | 23914 | 600 |
| IC-BIN [7] | 2 | – | 50K | – | 177 | 150 | 2176 | 1786 |
| LM [17] | 15 | – | 50K | – | 18273 | 3000 | 18273 | 3000 |
| RU-APC [50] | 14 | – | – | – | 5964 | 1380 | 5964 | 1380 |
| IC-MI [57] | 6 | – | – | – | 2067 | 300 | 5318 | 800 |
| TYO-L [24] | 21 | – | – | – | 1670 | 1670 | 1670 | 1670 |
| HOPE [58] | 28 | – | – | 50 | 188 | 188 | 3472 | 2898 |
@@ 3.1 Core datasets
BOP currently includes twelve datasets in a unified format. Sample test images are in Fig. 3 and dataset parameters in Tab. 1. Seven from the twelve were selected as core datasets: LM-O, T-LESS, ITODD, HB, YCB-V, TUD-L, and IC-BIN. Since 2019, methods must be evaluated on all of these core datasets to be considered for the main challenge awards (Sec. 5).
@@ 3.2 Training dataset for tasks on unseen objects
In 2023, as a training dataset for Tasks 4–6 (Sec. 2), we provided over 2M images in the BOP format showing more than 50K diverse objects (Fig. 3). The images were originally synthesized for MegaPose [33] using BlenderProc [5, 4, 6]. The objects are from the Google Scanned Objects [8] and ShapeNetCore [2] datasets. Note that symmetry transformations are not available for these objects, but could be identified as described in Sec. 2.3 of [25].
@@ 4.1 Experimental setup
Participants of the 2023 challenge were submitting results to the online evaluation system at bop.felk.cvut.cz from June 7, 2023 until the deadline on September 28, 2023. The evaluation scripts are publicly available in the BOP toolkit555github.com/thodan/bop_toolkit.
@@
Not a single pixel of test images may have been used for training and onboarding, nor the individual ground-truth annotations that are publicly available for test images of some datasets. Ranges of the azimuth and elevation camera angles, and a range of the camera-object distances determined by the ground-truth poses from test images are the only information about the test set that may have been used during training and onboarding. Only subsets of test images were used (see Tab. 1) to remove redundancies and speed up the evaluation, and only object instances for which at least of the projected surface area is visible were considered in the evaluation.
| # | Method | Year | Type | DNN per | Det./seg. | Refinement | Train im. | …type | Test im. | LM-O | T-LESS | TUD-L | IC-BIN | ITODD | HB | YCB-V | Time | |
| 1 | GPose2023 [61, 39] | 2023 | DNN | Object | Custom | Coord-guided | RGB-D | PBR+real | RGB-D | 79.4 | 91.4 | 96.4 | 73.7 | 70.4 | 95.0 | 92.8 | 85.6 | 2.67 |
| 2 | GPose2023-OfficialDet [61, 39] | 2023 | DNN | Object | Default GDRNPPDet | Coord-guided | RGB-D | PBR+real | RGB-D | 80.5 | 89.5 | 96.6 | 73.4 | 68.7 | 94.4 | 92.9 | 85.1 | 4.57 |
| 3 | GPose2023-PBR [61, 39] | 2023 | DNN | Object | Custom | Coord-guided | RGB-D | PBR+real | RGB-D | 79.4 | 89.0 | 93.1 | 73.7 | 70.4 | 95.0 | 90.1 | 84.4 | 2.86 |
| 4 | GDRNPP-PBRReal-RGBD-MModel [61, 39] | 2022 | DNN | Object | YOLOX | CIR | RGB-D | PBR+real | RGB-D | 77.5 | 87.4 | 96.6 | 72.2 | 67.9 | 92.6 | 92.1 | 83.7 | 6.26 |
| 5 | GDRNPP-PBR-RGBD-MModel [61, 39] | 2022 | DNN | Object | YOLOX | CIR | RGB-D | PBR | RGB-D | 77.5 | 85.2 | 92.9 | 72.2 | 67.9 | 92.6 | 90.6 | 82.7 | 6.26 |
| 6 | ZebraPoseSAT-EffnetB4-refined [53] | 2023 | DNN | Object | Default GDRNPPDet | CIR | RGB-D | PBR+real | RGB-D | 78.0 | 86.2 | 95.6 | 65.4 | 61.8 | 92.1 | 89.9 | 81.3 | 2.57 |
| 7 | GDRNPP-PBRReal-RGBD-MModel-Fast [61, 39] | 2022 | DNN | Object | YOLOX | Depth adjust. | RGB | PBR+real | RGB-D | 79.2 | 87.2 | 93.6 | 70.2 | 58.8 | 90.9 | 83.4 | 80.5 | 0.23 |
| 8 | OfficialDet-PFA-Mixpbr-RGB-D [27] | 2023 | DNN | Dataset | Default GDRNPPDet | PFA | RGB | PBR+real | RGB-D | 79.2 | 84.9 | 96.3 | 70.6 | 52.6 | 86.7 | 89.9 | 80.0 | 1.19 |
| 9 | GDRNPP-PBRReal-RGBD-MModel-Offi. [61, 39] | 2022 | DNN | Object | Default (synt+real) | CIR | RGB-D | PBR+real | RGB-D | 75.8 | 82.4 | 96.6 | 70.8 | 54.3 | 89.0 | 89.6 | 79.8 | 6.41 |
@@
| 11 | Extended_FCOS+PFA-MixPBR-RGBD [27] | 2022 | DNN | Dataset | Extended FCOS | PFA | RGB | PBR+real | RGB-D | 79.7 | 85.0 | 96.0 | 67.6 | 46.9 | 86.9 | 88.8 | 78.7 | 2.32 |
| 12 | Extended_FCOS+PFA-MixPBR-RGBD-Fast [27] | 2022 | DNN | Dataset | Extended FCOS | PFA | RGB | PBR+real | RGB-D | 79.2 | 77.9 | 95.8 | 67.1 | 46.0 | 86.0 | 88.0 | 77.1 | 0.64 |
| 13 | RCVPose3D-SingleModel-VIVO-PBR [63] | 2022 | DNN | Dataset | RCVPose3D | ICP | RGB-D | PBR+real | RGB-D | 72.9 | 70.8 | 96.6 | 73.3 | 53.6 | 86.3 | 84.3 | 76.8 | 1.34 |
| 14 | ZebraPoseSAT-EffnetB4+ICP(DefaultDet) [53] | 2022 | DNN | Object | Default (synt+real) | ICP | RGB | PBR+real | RGB-D | 75.2 | 72.7 | 94.8 | 65.2 | 52.7 | 88.3 | 86.6 | 76.5 | 0.50 |
| 15 | Extended_FCOS+PFA-PBR-RGBD [27] | 2022 | DNN | Dataset | Extended FCOS | PFA | RGB | PBR | RGB-D | 79.7 | 80.2 | 89.3 | 67.6 | 46.9 | 86.9 | 82.6 | 76.2 | 2.63 |
| 16 | SurfEmb-PBR-RGBD [15] | 2022 | DNN | Dataset | Default (PBR) | Custom | RGB-D | PBR | RGB-D | 76.0 | 82.8 | 85.4 | 65.9 | 53.8 | 86.6 | 79.9 | 75.8 | 9.05 |
| 17 | ZebraPoseSAT-EffnetB4 [53] | 2023 | DNN | Object | Default GDRNPPDet | – | RGB | PBR+real | RGB | 72.9 | 82.1 | 85.0 | 59.2 | 50.4 | 92.2 | 82.8 | 74.9 | 2.50 |
| 18 | GDRNPP-PBRReal-RGBD-SModel [61, 39] | 2022 | DNN | Dataset | YOLOX | Depth adjust. | RGB | PBR+real | RGB-D | 75.7 | 85.6 | 90.6 | 68.0 | 35.6 | 86.4 | 81.7 | 74.8 | 0.56 |
@@
| 20 | Coupled Iterative Refinement (CIR) [37] | 2022 | DNN | Dataset | Default (synt+real) | CIR | RGB-D | PBR+real | RGB-D | 73.4 | 77.6 | 96.8 | 67.6 | 38.1 | 75.7 | 89.3 | 74.1 | – |
| 21 | GPose2023-RGB [61, 39] | 2023 | DNN | Object | GDet2023 | CIR | RGB | PBR | RGB | 69.9 | 79.9 | 83.1 | 62.6 | 46.0 | 87.6 | 80.9 | 72.9 | 0.24 |
| 22 | GDRNPP-PBRReal-RGB-MModel[61, 39] | 2022 | DNN | Object | YOLOX | – | RGB | PBR+real | RGB | 71.3 | 78.6 | 83.1 | 62.3 | 44.8 | 86.9 | 82.5 | 72.8 | 0.23 |
| 23 | ZebraPoseSAT-EffnetB4 [53] | 2022 | DNN | Object | FCOS | – | RGB | PBR+real | RGB | 72.1 | 80.6 | 85.0 | 54.5 | 41.0 | 88.2 | 83.0 | 72.0 | 0.25 |
| 24 | ZebraPoseSAT-EffnetB4(DefaultDet) [53] | 2022 | DNN | Object | Default (synt+real) | – | RGB | PBR+real | RGB | 70.7 | 76.8 | 84.9 | 59.7 | 41.7 | 88.7 | 81.6 | 72.0 | 0.25 |
| 25 | ZebraPoseSAT-EffnetB4(PBR-DefaultDet) [53] | 2023 | DNN | Object | Default GDRNPPDet | – | RGB | PBR+real | RGB | 72.9 | 81.1 | 75.6 | 59.2 | 50.4 | 92.1 | 72.9 | 72.0 | 0.25 |
| 26 | ZebraPose-SAT [53] | 2022 | DNN | Object | FCOS | – | RGB | PBR+real | RGB | 72.1 | 78.7 | 86.1 | 54.9 | 37.9 | 84.7 | 82.8 | 71.0 | – |
| 27 | Extended_FCOS+PFA-MixPBR-RGB [27] | 2022 | DNN | Dataset | Extended FCOS | PFA | RGB | PBR+real | RGB | 74.5 | 77.8 | 83.9 | 60.0 | 35.3 | 84.1 | 80.6 | 70.9 | 3.02 |
| 28 | GDRNPP-PBR-RGB-MModel [61, 39] | 2022 | DNN | Object | YOLOX | – | RGB | PBR | RGB | 71.3 | 79.6 | 75.2 | 62.3 | 44.8 | 86.9 | 71.3 | 70.2 | 0.28 |
@@
| 31 | MRPE-PBRReal-RGB-SModel | 2023 | DNN | Dataset | Default GDRNPPDet | Render & com. | RGB | PBR+real | RGB-D | 74.4 | 75.8 | 82.4 | 55.0 | 36.8 | 77.0 | 84.3 | 69.4 | 0.10 |
| 32 | GDRNPP-PBRReal-RGB-SModel | 2022 | DNN | Dataset | YOLOX | CIR | RGB | PBR+real | RGB | 68.6 | 77.6 | 82.7 | 61.7 | 26.0 | 80.9 | 76.8 | 67.8 | 0.46 |